A Store-It-All storage facility requires users to enter a four-digit code, using numbers 0-9, to get through the gate. You need to get a box of winter clothes from your storage container, but you forgot the code. 

  1. How many possible codes are there if the digits cannot be repeated? 

There are Answer____ possible codes. 


1
Expert's answer
2022-01-25T14:54:11-0500

For the first place of the code, there are 10 different digits possible.

For the second place of the code, there are 9 different digits possible.

In the third place of the code, there are 8 different digits possible and,

in the fourth place of the code, there are 7 different digits possible.

Therefore, there are "10\times 9\times 8\times 7=5040" possible codes.